### v1

@@ 5 Understanding CHSH inequality with cards
As shown in Fig. 1, the objects to be measured are playing cards, characterized by two dichotomous properties: the color of their back (blue/red) and that of their front (black for spades and clubs, red for hearts and diamonds). As the system is classic, both properties can be measured for each card, the outcome of each measurement being ±$\pm 1$ (blu $=-1$, red $=1$; spades and clubs $=-1$, hearts and diamonds $=1$). The dealer (Charlie) prepares two decks of $N$ cards and gives the players (Alice and Bob) one deck of cards each. In this way, the two decks can be seen as formed by $N$ pairs of cards, characterized by two properties whose correlation we are interested in establishing and quantifying. Each player starts to measure the cards, being careful to maintain the given original order, that is, to preserve the possible correlation within each pair, and compiles a table as shown in Fig. 3, where the outcomes of all measurements are registered. It is important here to notice that being the system classical, we are able to measure all the quantities at the same time and calculate the value of $S_{i}$ with $i=1,..N$, for each row – i.e. each pair of cards. Though this could be considered a gimmick of the game - as in the real experiment only the average $\langle S\rangle$ can be calculated and has a precise statistical meaning – it allows us to easily convince everybody of the validity of CHSH inequality. Indeed, it is easy to recognize that for each row , whatever the colors of the card pair (see for instance the examples reported in Fig. 1), $S_{i}$ can assume only the values ±$\pm 2$, this meaning that its mean value $\langle S\rangle={\sum_{i}S_{i}}/{N}$ must be bound between ±$\pm 2$ as well, i.e. $|\langle S\rangle|\leq 2$.
This simple card game provides an elementary way to show CHSH inequality, accessible to anyone who knows how to calculate mean values. Moreover, it allows to naturally introduce the idea of perfect correlation ($\langle S\rangle=2$) between the properties of the two ensembles (when the cards of each pair are identical, as well as perfect anticorrelation ($\langle S\rangle=-2$), when all the cards in each pair have opposite values) and all the cases in between (as for instance $\langle S\rangle=0$ when no correlation at all is present). In presenting this game, a few crucial aspects should be highlighted and discussed:
@@
    A crucial difference between the classical CHSH game and the quantum one is that for the latter it is not possible to jointly measure all properties, i.e. $S_{i}$ cannot be evaluated. This difference should be clearly stated, making it explicit that calculating $S_{i}$ for each pair was a gimmick of the game to simplify the math. In a true experiment (either classical or quantum), the value of $\langle S\rangle$ is determined summing all the correlation functions by calculating the average of the products of the results of all pairs of measurements, i.e. by averaging each column of the table, rather than averaging over $S_{i}$. As long as the number of measurements is large enough (i.e. according to the Law of Large Numbers), the two procedures are equivalent and should lead to the same result. This emphasizes the statistical nature of the experimental result, clarifying the importance of collecting a large number of data, and of avoiding any possible bias in the choice of the measured quantities.
@@ 6 Staging the Nobel Prize-winning experiments
Having people acting the different phases of the experiment, even if only in a staged version of it, allows them to better understand the fundamental conceptual steps. Then, the real experiment can be discussed, both exploiting computer simulations, such as e.g. the one available from the Flytrap project [30, 31], and presenting one historical experiment, such as the one of the Nobel Prize Alain Aspect [32]. Flytrap simulation of the experiment allows performing step-by step measurements and verifying the violation of BI (when statistics is high enough).
As to Aspect’s experiment, performed in the early 1980s, the key points of the experimental setup are described. First of all, pairs of polarization-entangled photons are generated by a cascaded emission process in Calcium atoms, then the photons propagate to Alice’s and Bob’s laboratories where their polarization is measured. One of the key points of Aspect’s setup is that Alice and Bob randomly choose the polarization measurement basis – i.e. the orientation angle of the polarimeters – during photons’ time of flight: this is to avoid any possible influence between the measurements in the two laboratories. Aspect’s results for the values of $\langle S\rangle$ demonstrate a violation of the classical limit ±$\pm 2$ by 50 standard deviations. Overall, the experiments conducted by the Nobel Prize awardees have proven two key points:
@@ 7 Path implementation and preliminary validation
  • Public outreach events through the Italian Quantum Weeks initiative [8]
@@
  • Interactive elements requiring active participant involvement
@@
As probed by informal inquiry, most of the pupils in their 12th year had no previous knowledge on the subject. In Fig. 4 (a), we report, as a cloud of words, the answers to the preliminary question “which concepts and words do you associate with quantum mechanics?”. It is interesting to briefly comment on the most cited words and concepts. Among 55 answers, only three were structured sentences (“Branch of modern physics that deals with the study of matter and waves at the size of the atomic and subatomic order”, “Double slit experiment, atom structure, orbitals, interactions between atoms, string theory, subatomics, quanta, waves” but also “not applicable to reality.”), ten just provided scientist’ names (Bohr, Einstein, Heisenberg, Schroedinger and Planck, but also Maxwell and Huygens!). As shown in Fig. 4 (b), the most cited words were ‘numbers’ (most likely referring to quantum numbers), ‘waves’ and ‘particles’, followed by ‘quanta’ and ‘probability’. It is interesting to note that ‘relativity’ is also indicated several times. ‘Entanglement’ is mentioned twice and EPR once. When asked, only one pupil was able to tell what he knew about entanglement: “two entangles particles - even when farthest apart – remain in relation to each other”.
@@ I. Alice e Bob measure all polarisation direction of each photon $\times$
To gather information on the path effectiveness, at the end of the stage, pupils’ understanding was probed with a post-test, consisting both in open and multiple-choice items. Open questions regarded (a) the definition of entanglement, (b) the meaning of the hidden variables proposed by Einstein, and (c) the difference between correlation and causation. We recognize that answers to (a) are quite difficult to evaluate: while none of the pupils was able to provide a correct and complete definition of entanglement, most of them provided at least partially correct answers. As there is significant arbitrariness in how these answers can be evaluated, we decided not to further pursue this analysis (we can provide the original answers to the interested reader upon request). On the other hand, as far as (b) is concerned, most answers (17/26) capture the key point, that is the fact that hidden variables would provide a deterministic explanation of the probabilistic nature of QP, (7/26) answers were considered incorrect, as they mainly focus on the fact that these variables are unknown, while (2/26) pupils did not answer. It is interesting to note that question (c), was answered correctly only by (10/26) pupils. Among them, only one pupils provided a novel example, different from the one proposed during the class. Among wrong answers, most of them (8/26) confuse “causalità” (causality) with “casualità”, randomness , while (4/26) confuse the idea of spurious correlation with that of correlation without causation. This result shows that this key concept is not trivial, and that special care in the use of the similar terms (causalità/casualità) should be always taken. Concerning the multiple-choice part of the test, pupils were asked to chose among several statements (Q1) the ones that correctly describe BI and (Q2) the statements which describe the fundamental phases of the CHSH experiments, as enlisted in Table 1.
@@
As detailed in the caption of Fig. 5, almost 80% of the pupils chose the correct sentences regarding Charlie, Alice and Bob roles (sentences A, D, F), the random choice of Alice and Bob measurements (H and P), and that the observables measured by A and B are not commutative. Moreover, more than 60% of the pupils selected all the correct answers or made only one mistake, while less than 20% provided no answer or made several mistakes.
Finally, participants were asked (Q3) to chronologically order the selected phases. The results of this assessment component are summarized in Fig. 5. For Q3, we accepted two correct sequences:
@@
  • Meaningful learning outcomes can be achieved through targeted pedagogical approaches
This interpretation is supported by participant feedback shown in Fig. 6, which demonstrates strong overall satisfaction with the program. Particularly noteworthy is the correlation between assessment performance and positive survey responses, indicating both cognitive gains and affective engagement.
@@ 8 Conclusion
In recent years, with the emergence of quantum technologies, initiatives aimed at disseminating awareness about the foundations of Quantum Mechanics to the general public have become increasingly numerous. In this context, entanglement and its consequences are frequently used primarily to capture the public imagination, often waiving scientific rigor and sometimes leading to profound misconceptions. In fact, we believe that sharing the story of entanglement, the EPR paradox, and BI is essential, as these concepts are now woven into our cultural heritage and exemplify a landmark of scientific progress—one of the most remarkable achievements of twentieth-century physics. Furthermore, retracing the journey to experimentally verify BI violations over decades of precise, delicate work offers invaluable insights into the scientific process itself.
The real challenge lies in finding a way to explain the core character of this rather unique physical situation, so different from everyday experience, while avoiding misleading messages and misconceptions. Our experience has demonstrated that it is indeed possible to communicate these complex ideas to non-experts with only minimal use of mathematics, focusing on key elements like the structure of the entangled state and the derivation of BI. While a mathematical foundation is crucial, it must be paired with accessible, engaging communication that employs analogies and actively involves the audience. Integrating ‘gaming’ activities proved to be a crucial factor in the success of our educational approach, fostering curiosity and a hands-on understanding that brought these abstract concepts to life. In particular, any educational path regarding entanglement and EPR should address the issue of effectively presenting the Nobel Prize-winning experiments that eventually provided evidence for BI violation. Reproducing the real experiments based on entangled photon sources, though in principle possible even with off-the-shelf apparatus, is nevertheless beyond the possibilities of most educational initiatives. We showed that our staging activity is a valuable and viable solution to this problem. Besides the ease of implementation, this approach introduces personal experience and elements of embodied cognition, and facilitates grasping and retaining the key meaning of the otherwise quite cumbersome and complex experimental steps.
In conclusion, we believe that our strategy, which combines the use of ad-hoc designed staging activities to convey reasoning, offers an effective and engaging way to communicate the profound implications of quantum entanglement and its experimental verification, making these concepts accessible to a broader audience.
@@ Acknowledgements
We extend our gratitude to Marco G. Genoni and Andrea Smirne for their contributions in the early stages of this work and for the insightful discussions that greatly enriched this project. This project has been supported by the University of Modena and Reggio Emilia, by the no-profit organization Comitato Quantum, and by the Department of Science and High Technology of Insubria University.

### v2

@@ 5 Understanding CHSH inequality with cards
As shown in Fig. 1, the objects to be measured are playing cards, characterized by two dichotomous properties: the color of their back (blue/red) and that of their front (black for spades and clubs, red for hearts and diamonds). As the system is classic, both properties can be measured for each card, the outcome of each measurement being ±$\pm 1$ (blu $=-1$, red $=1$; spades and clubs $=-1$, hearts and diamonds $=1$). The dealer (Charlie) prepares two decks of $N$ cards and gives the players (Alice and Bob) one deck of cards each. In this way, the two decks can be seen as formed by $N$ pairs of cards, characterized by two properties whose correlation we are interested in establishing and quantifying. Each player starts to measure the cards, being careful to maintain the given original order, that is, to preserve the possible correlation within each pair, and compiles a table as shown in Fig. 3, where the outcomes of all measurements are registered. It is important here to notice that being the system classical, we are able to measure all the quantities at the same time and calculate the value of $S_{i}$ with $i=1,..N$, for each row – i.e. each pair of cards. Though this could be considered a gimmick of the game - as in the real experiment only the average $\langle S\rangle$ can be evaluated and has a precise statistical meaning – it allows us to easily convince everybody of the validity of CHSH inequality. Indeed, it is easy to recognize that for each row, whatever the colors of the card pair (see for instance the examples reported in Fig. 1), $S_{i}$ can assume only the values ±$\pm 2$, this meaning that its mean value $\langle S\rangle={\sum_{i}S_{i}}/{N}$ must be bound between ±$\pm 2$ as well, i.e. $|\langle S\rangle|\leq 2$.
This simple card game provides an elementary way to show CHSH inequality, accessible to anyone who knows how to calculate mean values. Moreover, it allows to naturally introduce the idea of perfect correlation between the properties of the two ensembles ($\langle S\rangle=2$, when the cards of each pair are identical), as well as perfect anticorrelation ($\langle S\rangle=-2$, when all the cards in each pair have opposite values) and all the cases in between (as for instance $\langle S\rangle=0$ when no correlation at all is present). In presenting this game, a few crucial aspects should be highlighted and discussed:
@@
    A crucial difference between the classical CHSH game and the quantum one is that for the latter it is not possible to jointly measure all properties, i.e. $S_{i}$ cannot be evaluated. This difference should be clearly stated, making it explicit that calculating $S_{i}$ for each pair was a gimmick of the game to simplify the math. In a true experiment (either classical or quantum), the value of $\langle S\rangle$ is determined summing all the correlation functions by calculating the average of the products of the results of all pairs of measurements, i.e. by averaging each column of the table, rather than averaging over $S_{i}$. As long as the number of measurements is large enough (i.e. according to the Law of Large Numbers), the two procedures are equivalent and should lead to the same result. This emphasizes the statistical nature of the experimental result, clarifying the importance of collecting a large number of data, and of avoiding any possible bias in the choice of the quantities to be measured.
@@ 6 Staging the Nobel Prize-winning experiments
Having people acting the different phases of the experiment, even if only in a staged version of it, allows them to better understand the fundamental conceptual steps.
Then, the real experiment can be discussed, both exploiting computer simulations, such as e.g. the one available from the Flytrap project [32, 33], and presenting one historical experiment, such as the one of the Nobel Prize Alain Aspect [34]. Flytrap simulation of the experiment allows performing step-by step measurements and verifying the violation of BI (when statistics is high enough). As to Aspect’s experiment, performed in the early 1980s, the key points of the experimental setup are described. First of all, pairs of polarization-entangled photons are generated by a cascaded emission process in Calcium atoms, then the photons propagate to Alice’s and Bob’s laboratories where their polarization is measured. One of the key points of Aspect’s setup is that Alice and Bob randomly choose the polarization measurement basis – i.e. the orientation angle of the polarimeters – during photon time of flight: this is to avoid any possible influence between the measurements in the two laboratories. Aspect’s results for the values of $\langle S\rangle$ demonstrate a violation of the classical limit ±$\pm 2$ by 50 standard deviations. Overall, the experiments conducted by the Nobel Prize awardees have proven two key points:
@@ 7 Path implementation and preliminary validation
  • Public outreach events through the Italian Quantum Weeks initiative. [8]
@@
  • Interactive elements requiring active participant involvement.
@@
As probed by informal inquiry, most of the pupils in their 12th year had no previous knowledge on the subject. In Fig. 4 (a), we report, as a cloud of words, the answers to the preliminary question “which concepts and words do you associate with quantum mechanics?”. It is interesting to briefly comment on the most cited words and concepts. Among 55 answers, only three were structured sentences (“Branch of modern physics that deals with the study of matter and waves at the size of the atomic and subatomic order”, “Double slit experiment, atomic structure, orbitals, interactions between atoms, string theory, subatomics, quanta, waves” but also “not applicable to reality.”), ten just provided scientist’ names (Bohr, Einstein, Heisenberg, Schroedinger and Planck, but also Maxwell and Huygens!). As shown in Fig. 4 (b), the most cited words were ‘numbers’ (most likely referring to quantum numbers), ‘waves’ and ‘particles’, followed by ‘quanta’ and ‘probability’. It is interesting to note that ‘relativity’ is also indicated several times. ‘Entanglement’ is mentioned twice and EPR once. When asked, only one pupil was able to tell what he knew about entanglement: “two entangles particles - even when farthest apart – remain in relation to each other”.
@@ I. Alice e Bob measure all polarisation direction of each photon $\times$
To gather information on the path effectiveness, at the end of the stage, pupils’ understanding was probed with a post-test, consisting both in open and multiple-choice items. Open questions regarded (a) the definition of entanglement, (b) the meaning of the hidden variables proposed by Einstein, and (c) the difference between correlation and causation. We recognize that answers to (a) are quite difficult to evaluate: while none of the pupils was able to provide a correct and complete definition of entanglement, most of them provided at least partially correct answers. As there is significant arbitrariness in how these answers can be evaluated, we decided not to further pursue this analysis (we can provide the original answers to the interested reader upon request). On the other hand, as far as (b) is concerned, most answers (17/26) capture the key point, that is the fact that hidden variables would provide a deterministic explanation of the probabilistic nature of QP, (7/26) answers were considered incorrect, as they mainly focus on the fact that these variables are unknown, while (2/26) pupils did not answer. It is interesting to note that question (c), was answered correctly only by (10/26) pupils. Among them, only one pupils provided a novel example, different from the one proposed during the class. Among wrong answers, most of them (8/26) confuse causality (causalità) with randomness (casualità) , while (4/26) confuse the idea of spurious correlation with that of correlation without causation. This result shows that this key concept is not trivial, and that (at least in Italy) special care in the use of the similar terms (causalità/casualità) should be always taken. Concerning the multiple-choice part of the test, pupils were asked to chose among several statements the ones that correctly describe BI (Q1) and the statements which describe the fundamental phases of the CHSH experiments (Q2), as enlisted in Table 1 (more then one choice was possible for each question).
@@
Concerning question (Q2), as detailed in the caption of Fig. 5, almost 80% of the pupils chose the correct sentences regarding Charlie, Alice and Bob roles (sentences A, D, F), the random choice of Alice and Bob measurements (H and P), and that the observables measured by A and B are not commutative. Moreover, more than 60% of the pupils selected all the correct answers or made only one mistake, while less than 20% provided no answer or made several mistakes.
Finally, participants were asked (Q3) to chronologically order the phases they selected in (Q2). The results of this assessment component are summarized in Fig. 5. For Q3, we accepted two correct sequences:
@@
  • Meaningful learning outcomes can be achieved through targeted pedagogical approaches.
This interpretation is supported by the participant feedback results shown in Fig. 6, which demonstrates strong overall satisfaction with the program. Particularly noteworthy are the high scores in both assessment performance and positive survey responses, indicating both cognitive gains and affective engagement.
@@ 8 Conclusion
In recent years, with the emergence of quantum technologies, initiatives aimed at disseminating awareness about the foundations of Quantum Physics to the general public have become increasingly numerous. In this context, entanglement and its consequences are frequently used primarily to capture the public imagination, often waiving scientific rigor and sometimes leading to profound misconceptions. In fact, we believe that sharing the story of entanglement, the EPR paradox, and BI is essential, as these concepts are now woven into our cultural heritage and exemplify a landmark of scientific progress—one of the most remarkable achievements of twentieth-century physics. Furthermore, retracing the journey to experimentally verify BI violations over decades of precise, delicate work offers invaluable insights into the scientific process itself.
The real challenge lies in finding a way to explain the core character of this rather unique physical situation, so different from everyday experience, while avoiding misleading messages and misconceptions. Our experience has demonstrated that it is indeed possible to communicate these complex ideas to non-experts with only minimal use of mathematics, focusing on key elements like the structure of the entangled state and the derivation of BI. While a mathematical foundation is crucial, it must be paired with accessible, engaging communication that employs analogies and actively involves the audience. Integrating ‘gaming’ activities proved to be a crucial factor in the success of our educational approach, fostering curiosity and a hands-on understanding that brought these abstract concepts to life. In particular, any educational path regarding entanglement and EPR should address the issue of effectively presenting the Nobel Prize-winning experiments that eventually provided evidence for BI violation. Reproducing the real experiments based on entangled photon sources, though in principle possible even with off-the-shelf apparatuses, is nevertheless beyond the possibilities of most educational initiatives. We showed that our staging activity is a valuable and viable solution to this problem. Besides the ease of implementation, this approach introduces personal experience and elements of embodied cognition, facilitating grasping and retaining the key meaning of the otherwise quite cumbersome and complex experimental steps.
In conclusion, we believe that our strategy, which uses of ad-hoc designed staging activities to convey reasoning, offers an effective and engaging way to communicate the profound implications of quantum entanglement and its experimental verification, making these concepts accessible to a broader audience.
@@ Acknowledgements
We thank Marco G. Genoni and Andrea Smirne for their contributions in the early stages of this work and for the insightful discussions. VDR and MGAP also thank Simone Cavazzoni, Gaia Forghieri, Giovanni Ragazzi, and Paolo Bordone for useful and entertaining discussions. This project has been supported by the University of Modena and Reggio Emilia, by the no-profit organization Comitato Quantum, and by the Department of Science and High Technology of Insubria University.